# Fluctuations in random field Ising models

### v1

@@ Abstract
This paper establishes a CLT for linear statistics of the form $\langle\mathbf{q},\mathnormal{\bm{\sigma}}\rangle$ with quantitative Berry-Esseen bounds, where $\mathnormal{\bm{\sigma}}$ is an observation from an exponential family with a quadratic form as its sufficient statistic, in the “high-temperature” regime. We apply our general result to random field Ising models with both discrete and continuous spins. To demonstrate the generality of our techniques, we apply our results to derive both quenched and annealed CLTs in various examples, which include Ising models on some graph ensembles of common interest (Erdős-Rényi, regular, dense bipartite), and the Hopfield spin glass model. Our proofs rely on a combination of Stein’s method of exchangeable pairs and Chevet type concentration inequalities.
Berry-Esseen bounds,
Central limit theorem,
Hopfield model,
Mean-Field approximation,
Random field Ising Model,
@@ 1 Introduction
Let $\mathnormal{\bm{\sigma}}:=(\sigma_{1},\ldots,\sigma_{n})\in[-1,1]^{n}$ be a random vector drawn from the following quadratic interaction model with external field $\mathnormal{\mathbf{c}}:=(c_{1},\ldots,c_{n})\in\mathbb{R}^{n}$:
@@
Here, each $\mu_{i}$ is a base-measure supported on $[-1,1]$, and $\mathbf{A}_{n}$ is a symmetric matrix with zeros on the diagonal. We will refer to $\mathbf{c}$ as the field vector, and the factor $Z_{n}$ as the normalizing constant/partition function of model (1.1). Note that the inverse temperature parameter for typical Gibbs measures is absorbed into the coupling matrix $\mathbf{A}_{n}$. We omit the dependence of $(\mathbf{A}_{n},\mathbf{c},\{\mu_{i}\}_{1\leq i\leq n})$ on the measure $\operatorname{\mathbb{P}}$ and normalizing constant $Z_{n}$ for simplicity of notation.
@@
for $\mathbf{q}:=(q_{1},\ldots,q_{n})\in\mathbb{R}^{n}$, normalized to have $\|\mathbf{q}\|=1$, when the matrix $\mathbf{A}_{n}$ is in the so called high-temperature Mean-Field regime of statistical physics (see Assumptions 2.1 and 2.2 below). Our main motivation for studying limit theorems under (1.1) comes from the following example:
@@
Consider an observation from model (1.1) where the base measures $\mu_{i}=\mu$ are the same, and the field vector $\mathbf{c}=(c_{1},\ldots,c_{n})$ is a fixed realization of $n$ i.i.d. random variables from some probability distribution, say $F$. The most studied variant of this model is the random field Ising model (RFIM, sometimes referred to as the disordered Ising model), where $\mu_{i}$s are supported on $\{-1,1\}$ and $\mathbf{A}_{n}$ represents the (scaled) adjacency matrix of a graph. Two commonly studied graph ensembles are the complete graph (see [34, 12, 20, 22]), and a neighborhood graph which is a subgraph of the $d$-dimensional lattice (see [16, 9]). In contrast, we look at more general spin-glass models, where the entries of $\mathbf{A}_{n}$ can have both positive and negative entries, and also allow for general compactly supported probability measures $\mu$ (discrete or continuous). Do note that our results apply only in the so called high-temperature regime (see 2.1).
@@ 1.1 Main Contributions
    In Section 2.3, we provide our main results regarding concentration and fluctuations of the linear statistic $T_{n}=\mathbf{q}^{\top}\bm{\sigma}$, where $\mathnormal{\bm{\sigma}}$ is a random sample from the distribution $\operatorname{\mathbb{P}}$ (see (1.1)). This includes the following: (i) Law of large numbers (LLN) and concentration bounds (in 2.3), (ii) a CLT with finite sample Berry-Esseen bounds with a somewhat implicit centering (in 2.4), and (iii) another CLT with Berry-Esseen bounds with a more explicit centering, under mildly stronger assumptions (see 2.5). The LLN is stated under minimal assumptions, where $\mathbf{q}$ is an arbitrary vector. For the Berry-Esseen bounds to converge to 00, we additionally require $\mathbf{q}$ to be an approximate delocalized eigenvector of the coupling matrix $\mathbf{A}_{n}$. In fact, the corresponding approximate eigenvalue shows up in the limiting normal distribution. We note that as our results are non-asymptotic with explicit error bounds, they can be readily applied to provide quenched and annealed limits for $T_{n}$ in a variety of quadratic interaction models, as illustrated in 2.6.
@@
    Next, in Section 2.4, we apply our results in the context of the random field Ising model, introduced in 1.1. To the best of our knowledge, only the average magnetization of the Curie-Weiss (complete graph) RFIM has been understood in the literature [12]. In 2.6, we state our main result for RFIMs with general coupling matrices $\mathbf{A}_{n}$, which include adjacency matrices of several graph ensembles (Erdős-Rényi graphs, regular graphs, complete bipartite graph), as well as the celebrated spin-glass Hopfield model [21]. Moving beyond complete graphs brings in new technical challenges which we overcome with new proof techniques that involve Stein’s method of exchangeable pairs [10, 35], new Chevet type inequalities (see Lemma 4.1), and new contraction bounds on certain “local” averages (see (1.5) and Lemma 3.2).
@@ 1.2 Notation
For two measures $\kappa_{1},\kappa_{2}$ on the same probability space, define the Kullback-Leibler divergence between $\kappa_{1}$ and $\kappa_{2}$ as
$\mathrm{KL}(\kappa_{1}|\kappa_{2}):=\begin{cases}\operatorname{\mathbb{E}}_{\kappa_{1}}\log{\frac{d\kappa_{1}}{d\kappa_{2}}}&\mbox{if }\kappa_{1}\ll\kappa_{2},\\ \infty&\mbox{otherwise}.\end{cases}$
Given an arbitrary symmetric $p\times p$ real valued matrix $\mathbf{M}_{n}$, let $\lVert\mathbf{M}_{n}\rVert$, $\lVert\mathbf{M}_{n}\rVert_{r}$, and $\lVert\mathbf{M}_{n}\rVert_{F}$ denote the $(2,2)$-operator norm, the $(r,r)$-operator norm for $r\in(2,\infty]$, and the Frobenius norm of $\mathbf{M}_{n}$ respectively. We also let $\mathrm{Off}(\mathbf{M}_{n})$ denote the off-diagonal operator that sets the diagonal entries in $\mathbf{M}_{n}$ to 0. For a vector $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$, let $\lVert\mathbf{a}\rVert$ and $\lVert\mathbf{a}\rVert_{r}$ denote the $\ell_{2}$ and $\ell_{r}$-vector norms, for $r\in(2,\infty]$. Let $\mbox{Diag}(\mathbf{a})$ denote the $p\times p$ diagonal matrix with diagonal entries $a_{1},\ldots,a_{n}$. We denote the $n$ dimensional all-zero vector, all-one vector, and $n\times n$ identity matrix by $\mathbf{0}$, $\mathbf{1}$, and $\mathbf{I}_{n}$ respectively. For $x\in\mathbb{R}$, $\lceil x\rceil$ will denote the smallest integer larger than or equal to $x$. With $\mathbf{A}_{n}$ as in (1.1), define
$\alpha_{n}:=\max_{1\leq i\leq p}\sum_{j=1}^{n}\mathbf{A}_{n}^{2}(i,j).$ (1.4)
[…]
@@
Unless otherwise mentioned, all probabilities are computed under $\operatorname{\mathbb{P}}$. Finally, we introduce the notion of exponential tilting for probability measures below.
@@
Consider the $\mu_{i}$s in (1.1), for $1\leq i\leq n$. For $\theta\in\mathbb{R}$, define the tilted measures $\mu_{i,\theta}$ as follows:
@@
[…]
$\psi_{i}^{\prime}(\theta)=\operatorname{\mathbb{E}}_{\mu_{i,\theta}}[Z],\quad\psi_{i}^{\prime\prime}(\theta)={\rm Var}_{\mu_{i,\theta}}[Z].$
@@ 2.1 Assumptions and Mean-Field approximation
Note that the SHT assumption in (2.3) is commonly referred to as the high-temperature Dobrushin uniqueness condition in the Gibbs distribution literature; see e.g. [18, Equation 1.1] and [25, Equation 2.2]. Different forms of such high-temperature assumptions on $\mathbf{A}_{n}$ have gained significant popularity in the recent years (see e.g., [29, 2, 31, 28, 7]). The following matrix lemma provides an ordering between the three high-temperature regimes stated above.
@@
Next, we introduce the Mean-Field approximation of $\operatorname{\mathbb{P}}$ (see (1.1)), which will play an instrumental role in describing the asymptotic distribution of $T_{n}$. Recall the definition of the normalizing constant $Z_{n}$ from (1.2). The Gibbs variational principle implies
$\log{Z_{n}}=\sup_{Q}\left(\operatorname{\mathbb{E}}_{Q}\left[\frac{1}{2}\operatorname{\mathnormal{\bm{\sigma}}}^{\top}\mathbf{A}_{n}\operatorname{\mathnormal{\bm{\sigma}}}+\mathnormal{\mathbf{c}}^{\top}\operatorname{\mathnormal{\bm{\sigma}}}\right]-\mathrm{KL}\big{(}Q|\prod_{i=1}^{n}\mu_{i}\big{)}\right),$
where the supremum is taken over all probability measures $Q$ on $[-1,1]^{n}$ (see [37]), and is attained at $Q=\operatorname{\mathbb{P}}$. The Mean-Field approximation (see [5, 6]) for $\log Z_{n}$ arises by restricting the above supremum to the family of product measures, i.e.,
$\log{Z_{n}}$ $\geq\sup_{Q=\prod_{i=1}^{n}Q_{i}}\left(\operatorname{\mathbb{E}}_{Q}\left[\frac{1}{2}\operatorname{\mathnormal{\bm{\sigma}}}^{\top}\mathbf{A}_{n}\operatorname{\mathnormal{\bm{\sigma}}}+\mathnormal{\mathbf{c}}^{\top}\operatorname{\mathnormal{\bm{\sigma}}}\right]-\mathrm{KL}\big{(}Q|\prod_{i=1}^{n}\mu_{i}\big{)}\right)=\sup_{\mathbf{z}\in[-1,1]^{n}}F(\mathbf{z}),$ (2.4)
@@
The equality (2.4) follows from simplifying the supremum on the middle term by optimizing over product measures $Q=\prod_{i=1}^{n}Q_{i}$ where each $Q_{i}$ has mean $z_{i}$; see e.g. Proposition 1 in [38]. The following lemma demonstrates that $F(\cdot)$ has a unique maximizer under a high-temperature assumption.
@@
$\frac{dQ^{\mathrm{prod}}}{d\prod_{i=1}^{n}\mu_{i}}(\operatorname{\mathnormal{\bm{\sigma}}}):\propto\prod_{i=1}^{n}\exp(\phi_{i}(u_{i})\sigma_{i})$ (2.7)
corresponds to the maximizer in the first supremum of (2.4), as the mean of the $i$th marginal of $Q^{\mathrm{prod}}$ is exactly $u_{i}$.
@@
The Mean-Field approximation $Q^{\mathrm{prod}}=\prod_{i=1}^{n}Q^{\mathrm{prod}}_{i}$ motivates the natural centered statistic
$T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}):=T_{n}-\sum_{i=1}^{n}q_{i}\operatorname{\mathbb{E}}_{Q^{\mathrm{prod}}_{i}}[\sigma_{i}]=\sum_{i=1}^{n}q_{i}(\sigma_{i}-u_{i}).$ (2.8)
@@
We refer to the above condition as Strong Mean-Field assumption, as this implies the conclusion $\lVert\mathbf{A}_{n}\rVert_{F}^{2}=o(n)$, which has been used as a Mean-Field assumption in recent literature (see [3, 38, 29, 13, 14]). We motivate our stronger assumption by presenting our first theorem.
@@ 2.2 Weak laws and concentration inequalities
Further, if the SHT condition (2.3) holds, then there exists positive constants $h_{1}$ and $h_{2}$, such that given any $t>0$, we have:
@@
This is what one would expect if $\operatorname{\mathbb{P}}$ were a product distribution with mean vector ${\mathbf{u}}$, so this bound is the best possible, and one cannot expect a better result even under independence. Secondly, 2.3 gives non-trivial conclusions even without the SMF condition (2.9). To see this, take ${\bf q}=n^{-1/2}{\bf 1}_{n}$. The law of large numbers
@@
holds as soon as $\alpha_{n}\to 0$. It is well known that Mean-Field centering can fail if $\alpha_{n}$ stays bounded away from 00 (see e.g., [32, 7]), so the requirement $\alpha_{n}\to 0$ in the above display is tight.
@@ 2.3 Central Limit Theorems
Suppose that $\operatorname{\mathnormal{\bm{\sigma}}}\sim\operatorname{\mathbb{P}}$ and the MHT condition (2.2) holds. Let $\mathbf{u}$ be the Mean-Field optimizers as in Lemma 2.2. For all $i=1,\ldots,n$, let $\nu_{i}:=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)\psi_{j}^{\prime}(c_{j})$. Set $\upsilon_{n}:=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\psi_{i}^{\prime\prime}(c_{i})>0$, and define the following:
@@
Fix $\lambda_{n}\in\mathbb{R}$, and set $\bm{\epsilon}:=\mathbf{A}_{n}\mathbf{q}-\lambda_{n}\mathbf{q}$. Suppose $W_{n}\sim\operatorname{N}\left(0,\frac{\upsilon_{n}}{1-\lambda_{n}\upsilon_{n}}\right)$. We then have
$\;\;\;d_{KS}\left(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})\,,\,W_{n}\right)\lesssim\sqrt{R_{1n}}+\sqrt{\operatorname{\alpha_{\mathnormal{n}}}R_{2n}}+\frac{\sqrt{R_{3n}}+\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}+\lVert\mathbf{q}\rVert_{\infty}}{\upsilon_{n}}+\lVert\bm{\epsilon}\rVert_{2}.$ (2.13)
Here the implicit constant only depends on $\rho$ (see (2.2)).
@@
[…]
@@
Note that, for $T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})$ to converge weakly to a mean 00 Gaussian distribution, the SMF condition (2.9) is required. If $\sqrt{n}\alpha_{n}$ converges to a positive real instead, then the limiting Normal may not be centered at 00 for specific choices of $\mathbf{A}_{n}$ and $\mathbf{c}$; see e.g. [14, Example 1.3]. In this sense, the SMF condition is tight.
For both Theorems 2.3 and 2.4, we have centered $T_{n}(\operatorname{\mathnormal{\bm{\sigma}}})$ by subtracting $\mathbf{q}^{\top}\mathbf{u}$, where the Mean-Field solution $\mathbf{u}$ is obtained from Lemma 2.2. Typically, ${\mathbf{u}}$ needs to be computed using numerical optimization methods. The purpose of the following result is to show that, at the expense of some additional error terms, $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}\sigma_{i}$ has Gaussian fluctuations under $\operatorname{\mathbb{P}}$, around a more explicit function of $\mathbf{q}$, $\mathbf{A}_{n}$, and $\mathbf{c}$.
@@
Suppose $\operatorname{\mathnormal{\bm{\sigma}}}\sim\operatorname{\mathbb{P}}$ and consider the same notation as in 2.4. Suppose that the MHT condition (2.2) holds (recall that $\bm{\epsilon}=\mathbf{A}_{n}\mathbf{q}-\lambda_{n}\mathbf{q}$). Set $\Psi^{\prime}(\mathbf{c}):=\left(\psi_{1}^{\prime}(c_{1}),\ldots,\psi_{n}^{\prime}(c_{n})\right)$ and
@@
    $\lesssim\sqrt{R_{2n}}\left(\sqrt{R_{1n}}+\sqrt{\alpha_{n}}+\lVert\bm{\epsilon}\rVert\right)+\sqrt{R_{1n}}+\frac{\sqrt{R_{3n}}+\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}+\lVert\mathbf{q}\rVert_{\infty}}{\upsilon_{n}}+R_{4n}+\big{|}\bm{\epsilon}^{\top}\Psi^{\prime}(\mathbf{c})\big{|}.$ (2.15)
    In both cases, the implicit constant depends only on $\rho$ from (2.2).

### v2

@@ Abstract
This paper establishes a CLT for linear statistics of the form $\langle\mathbf{q},\mathnormal{\bm{\sigma}}\rangle$ with quantitative Berry-Esseen bounds, where $\mathnormal{\bm{\sigma}}$ is an observation from an exponential family with a quadratic form as its sufficient statistic, in the “high-temperature” regime. We apply our general result to random field Ising models with both discrete and continuous spins. To demonstrate the generality of our techniques, we apply our results to derive both quenched and annealed CLTs in various examples, which include Ising models on some graph ensembles of common interest (Erdős-Rényi, regular, random graphons), and the Hopfield spin glass model. Our proofs rely on a combination of Stein’s method of exchangeable pairs and Chevet-type concentration inequalities.
@@ 1 Introduction
Let $\mathnormal{\bm{\sigma}}:=(\sigma_{1},\ldots,\sigma_{n})^{\top}\in[-1,1]^{n}$ be a random vector drawn from the following quadratic interaction model with external field $\mathnormal{\mathbf{c}}:=(c_{1},\ldots,c_{n})^{\top}\in\mathbb{R}^{n}$:
@@
[…]
@@
for $\mathbf{q}:=(q_{1},\ldots,q_{n})^{\top}\in\mathbb{R}^{n}$, normalized to have $\|\mathbf{q}\|=1$, when the matrix $\mathbf{A}_{n}$ is in the so called high-temperature Mean-Field regime of statistical physics (see Assumptions 2.1 and 2.2 below). Our main motivation for studying limit theorems under (1.1) comes from the following example:
@@
Consider an observation from model (1.1) where the base measures $\mu_{i}=\mu$ are the same, and the field vector $\mathbf{c}=(c_{1},\ldots,c_{n})^{\top}$ is a fixed realization of $n$ i.i.d. random variables from some probability distribution, say $F$. The most studied variant of this model is the random field Ising model (RFIM, sometimes referred to as the disordered Ising model), where $\mu_{i}$s are supported on $\{-1,1\}$ and $\mathbf{A}_{n}$ represents the (scaled) adjacency matrix of a graph. Two commonly studied graph ensembles are the complete graph (see [44, 18, 30, 32]), and a neighborhood graph which is a subgraph of the $d$-dimensional lattice (see [14, 13, 25, 15, 24, 23, 22]). In contrast, we look at more general spin-glass models, where the entries of $\mathbf{A}_{n}$ can have both positive and negative entries, and also allow for general compactly supported probability measures $\mu$ (discrete or continuous). Do note that our results apply only in the so called high-temperature regime (see 2.1).
@@ 1.1 Main Contributions
[…]
@@
    Next, in Section 2.5, we apply our results in the context of the random field Ising model, introduced in 1.1. To the best of our knowledge, only the average magnetization of the Curie-Weiss (complete graph) RFIM has been understood in the literature [18]. In 2.6, we state our main result for RFIMs with general coupling matrices $\mathbf{A}_{n}$, which include adjacency matrices of several graph ensembles (Erdős-Rényi graphs, regular graphs, random graphons), as well as the celebrated spin-glass Hopfield model [31]. Moving beyond complete graphs brings in new technical challenges which we overcome with new proof techniques that involve Stein’s method of exchangeable pairs [16, 45], new Chevet-type inequalities (see Lemma 4.1), and new contraction bounds on certain “local” averages (see (1.5) and Lemma 3.2).
@@ 1.2 Notation
For two measures $\pi_{1},\pi_{2}$ on the same probability space, define the Kullback-Leibler divergence between $\pi_{1}$ and $\pi_{2}$ as
$\mathrm{KL}(\pi_{1}|\pi_{2}):=\begin{cases}\operatorname{\mathbb{E}}_{\pi_{1}}\log{\frac{d\pi_{1}}{d\pi_{2}}}&\mbox{if }\pi_{1}\ll\pi_{2},\\ \infty&\mbox{otherwise}.\end{cases}$
Given an arbitrary symmetric $n\times n$ real valued matrix $\mathbf{M}_{n}$, let $\lVert\mathbf{M}_{n}\rVert$, $\lVert\mathbf{M}_{n}\rVert_{r}$, and $\lVert\mathbf{M}_{n}\rVert_{F}$ denote the $(2,2)$-operator norm, the $(r,r)$-operator norm for $r\in(2,\infty]$, and the Frobenius norm of $\mathbf{M}_{n}$ respectively. We also let $\mathrm{Off}(\mathbf{M}_{n})$ denote the off-diagonal operator that sets the diagonal entries in $\mathbf{M}_{n}$ to 0. For a vector $\mathbf{a}=(a_{1},\ldots,a_{n})^{\top}\in\mathbb{R}^{n}$, let $\lVert\mathbf{a}\rVert$ and $\lVert\mathbf{a}\rVert_{r}$ denote the $\ell_{2}$ and $\ell_{r}$-vector norms, for $r\in(2,\infty]$. Let $\mbox{Diag}(\mathbf{a})$ denote the $n\times n$ diagonal matrix with diagonal entries $a_{1},\ldots,a_{n}$. We denote the $n$ dimensional all-zero vector, all-one vector, and $n\times n$ identity matrix by $\mathbf{0}$, $\mathbf{1}$, and $\mathbf{I}_{n}$ respectively. For $\in\mathbb{R}, x\lceil x\rceil$ will denote the smallest integer larger than or equal to $x$. With $\mathbf{A}_{n}$ as in (1.1), define
$\alpha_{n}:=\max_{1\leq i\leq n}\sum_{j=1}^{n}\mathbf{A}_{n}^{2}(i,j).$ (1.4)
For nonnegative sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we write $a_{n}\lesssim b_{n}$ or $a_{n}=O(b_{n})$, if there exists a constant $C>0$ free of $n$ such that $a_{n}\leq Cb_{n}$. We use $a_{n}\ll b_{n}$ or $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. We also use the standard $o_{P},O_{p}$ notations for a sequence of random variables. The notation $\overset{d}{\longrightarrow}$ denotes weak convergence/convergence in distribution, and $\overset{P}{\longrightarrow}$ denotes convergence in probability. For a random vector $\operatorname{\mathnormal{\bm{\sigma}}}$} from the model (1.1), we define the random vector of local fields as $\mathbf{m}:=(m_{1},\ldots,m_{n})^{\top}$ where
@@
[…]
@@
Consider the non-degenerate base measures $\mu_{i}$s in (1.1), for $1\leq i\leq n$. For $\theta\in\mathbb{R}$, define the tilted measures $\mu_{i,\theta}$ as follows:
@@
Let $\psi_{i}(\cdot)$ denote the log moment generating function of $\mu_{i}$, defined by $\psi_{i}(\theta):=\log\int e^{\theta z}d\mu_{i}(z)$ for $\theta\in\mathbb{R}$. Then, standard exponential family calculations gives
$\psi_{i}^{\prime}(\theta)=\operatorname{\mathbb{E}}_{Z\sim\mu_{i,\theta}}[Z],\quad\psi_{i}^{\prime\prime}(\theta)={\rm Var}_{Z\sim\mu_{i,\theta}}[Z].$
@@ 2.1 Assumptions and Mean-Field approximation
Note that the SHT assumption in (2.3) is commonly referred to as the high-temperature Dobrushin uniqueness condition in the Gibbs distribution literature; see e.g. [28, Equation 1.1] and [35, Equation 2.2]. Different forms of such high-temperature assumptions on $\mathbf{A}_{n}$ have gained significant popularity in the recent years (see e.g., [39, 2, 41, 38, 11]). The following matrix lemma provides an ordering between the three high-temperature regimes stated above.
@@
[…]
$\log{Z_{n}}=\sup_{\operatorname{\mathbb{Q}}}\left(\operatorname{\mathbb{E}}_{\operatorname{\mathbb{Q}}}\left[\frac{1}{2}\operatorname{\mathnormal{\bm{\sigma}}}^{\top}\mathbf{A}_{n}\operatorname{\mathnormal{\bm{\sigma}}}+\mathnormal{\mathbf{c}}^{\top}\operatorname{\mathnormal{\bm{\sigma}}}\right]-\mathrm{KL}\big(\operatorname{\mathbb{Q}}|\prod_{i=1}^{n}\mu_{i}\big)\right),$
[…]
$\log{Z_{n}}$ $\geq\sup_{\operatorname{\mathbb{Q}}=\prod_{i=1}^{n}\operatorname{\mathbb{Q}}_{i}}\left(\operatorname{\mathbb{E}}_{\operatorname{\mathbb{Q}}}\left[\frac{1}{2}\operatorname{\mathnormal{\bm{\sigma}}}^{\top}\mathbf{A}_{n}\operatorname{\mathnormal{\bm{\sigma}}}+\mathnormal{\mathbf{c}}^{\top}\operatorname{\mathnormal{\bm{\sigma}}}\right]-\mathrm{KL}\big(\operatorname{\mathbb{Q}}|\prod_{i=1}^{n}\mu_{i}\big)\right)=\sup_{\mathbf{z}\in[-1,1]^{n}}F(\mathbf{z}),$ (2.4)
@@
The equality (2.4) follows from simplifying the supremum on the middle term by optimizing over product measures $\operatorname{\mathbb{Q}}=\prod_{i=1}^{n}\operatorname{\mathbb{Q}}_{i}$ where each $\operatorname{\mathbb{Q}}_{i}$ has mean $z_{i}$; see e.g. Proposition 1 in [48]. The following lemma demonstrates that $F(\cdot)$ has a unique maximizer under a high-temperature assumption.
@@
$\frac{d\mathbb{Q}^{\mathrm{prod}}}{d\prod_{i=1}^{n}\mu_{i}}(\operatorname{\mathnormal{\bm{\sigma}}}):\propto\prod_{i=1}^{n}\exp(\phi_{i}(u_{i})\sigma_{i})$ (2.7)
corresponds to the maximizer in the first supremum of (2.4), as the mean of the $i$th marginal of $\mathbb{Q}^{\mathrm{prod}}$ is exactly $u_{i}$.
@@
The Mean-Field approximation $\mathbb{Q}^{\mathrm{prod}}=\prod_{i=1}^{n}\mathbb{Q}^{\mathrm{prod}}_{i}$ motivates the natural centered statistic
$T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}):=T_{n}-\sum_{i=1}^{n}q_{i}\operatorname{\mathbb{E}}_{\mathbb{Q}^{\mathrm{prod}}_{i}}[\sigma_{i}]=\sum_{i=1}^{n}q_{i}(\sigma_{i}-u_{i}).$ (2.8)
@@
We refer to the above condition as Strong Mean-Field assumption, as this implies the conclusion $\lVert\mathbf{A}_{n}\rVert_{F}^{2}=o(\sqrt{n})$, which is stronger than the more commonly used Mean-Field assumption $\lVert\mathbf{A}_{n}\rVert_{F}^{2}=o(n)$ (see [3, 48, 39, 19, 20]). We motivate our stronger assumption by presenting our first theorem.
@@ 2.2 Weak laws and concentration inequalities
Further, if the SHT condition (2.3) holds, then there exist positive constants $h_{1}$ and $h_{2}$, such that given any $t>0$, we have:
@@
[…]
@@
holds as soon as $\alpha_{n}\to 0$. It is well known that Mean-Field centering can fail if $\alpha_{n}$ stays bounded away from 0 (see e.g., [42, 11]), so the requirement $\alpha_{n}\to 0$ in the above display is tight.
@@ 2.3 Central Limit Theorems
Suppose that $\operatorname{\mathnormal{\bm{\sigma}}}\sim\operatorname{\mathbb{P}}$ and the MHT condition (2.2) holds. Let $\mathbf{u}$ be the Mean-Field optimizers as in Lemma 2.2. For all $i=1,\ldots,n$, let $\nu_{i}:=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)\psi_{j}^{\prime}(c_{j})$. Assume $\upsilon_{n}:=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\psi_{i}^{\prime\prime}(c_{i})\geq\kappa$ for some constant $\kappa\in(0,1)$, and define the following:
@@
Fix $\lambda_{n}$ such that $|\lambda_{n}|\leq\rho$ (see (2.2)), and set $\bm{\epsilon}:=\mathbf{A}_{n}\mathbf{q}-\lambda_{n}\mathbf{q}$. For $W_{n}\sim\operatorname{N}\left(0,\frac{\upsilon_{n}}{1-\lambda_{n}\upsilon_{n}}\right)$, we have
$\;\;\;d_{KS}\left(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})\,,\,W_{n}\right)\lesssim\sqrt{R_{1n}}+\sqrt{\operatorname{\alpha_{\mathnormal{n}}}R_{2n}}+\sqrt{R_{3n}}+\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}+\lVert\mathbf{q}\rVert_{\infty}+\lVert\bm{\epsilon}\rVert.$ (2.13)
Here the implicit constant only depends on $\rho$ and $\kappa$.
@@
$\limsup\limits_{n\to\infty}|\lambda_{n}|\leq\limsup\limits_{n\to\infty}\,\lVert\mathbf{A}_{n}\mathbf{q}\rVert\leq\limsup\limits_{n\to\infty}\,\lVert\mathbf{A}_{n}\rVert_{4}\leq\rho<1,$
@@
Note that, for $T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})$ to converge weakly to a mean 0 Gaussian distribution, the SMF condition (2.9) is required. If $\sqrt{n}\alpha_{n}$ converges to a positive real instead, then the limiting Normal may not be centered at 0 for specific choices of $\mathbf{A}_{n}$ and $\mathbf{c}$; see e.g. [20, Example 1.3]. In this sense, the SMF condition is tight.
[…]
@@
Suppose $\operatorname{\mathnormal{\bm{\sigma}}}\sim\operatorname{\mathbb{P}}$ and consider the same notation as in 2.4. Suppose that the MHT condition (2.2) holds and assume $|\lambda_{n}|\leq\rho,\upsilon_{n}\geq\kappa>0$. Recall that $\bm{\epsilon}=\mathbf{A}_{n}\mathbf{q}-\lambda_{n}\mathbf{q}$. Set $\Psi^{\prime}(\mathbf{c}):=\left(\psi_{1}^{\prime}(c_{1}),\ldots,\psi_{n}^{\prime}(c_{n})\right)^{\top}$ and
@@
    $\lesssim\sqrt{R_{2n}}\left(\sqrt{R_{1n}}+\sqrt{\alpha_{n}}+\lVert\bm{\epsilon}\rVert\right)+\sqrt{R_{1n}}+\sqrt{R_{3n}}+\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}+\lVert\mathbf{q}\rVert_{\infty}+R_{4n}+\big|\bm{\epsilon}^{\top}\Psi^{\prime}(\mathbf{c})\big|+\|\bm{\epsilon}\|.$ (2.15)
    In both cases, the implicit constant depends only on $\rho$ from (2.2) and $\kappa$.